OpenStudy (anonymous):

how to find the solution for differential equation of an RC Circuit when the equation is: RC(dUc/dt)+Uc=E

OpenStudy (vincent-lyon.fr):

There are different ways. Either you already know the general solution to such a first order linear equation, or you use separation of variables. Put Uc on one side and time on the other side: \(\Large \frac{dU_c}{dt} =\frac {E-U_c}{RC} \) , then \(\Large \frac{dU_c}{U_c-E} =-\frac {dt}{RC} \) Integrate both sides from initial conditions (for instance t=0 and Uc=0) to any time and voltage t and Uc. This develops to: \(\ln \Large \frac{U_c-E}{0-E} =-\frac {t-0}{RC} \) Take exponential and rearrange to: \(U_c =E\: (1-\Large e\;\Large ^{ \frac {-t}{RC}}) \)

OpenStudy (anonymous):

Thank's :) btw do we use the same method to find it for RL & RLC??

OpenStudy (vincent-lyon.fr):

For RL you can, but RLC needs help of mathematicians because it is a second order differential equation.

OpenStudy (anonymous):

ouh ok thx for your help :)